OpenStudy (anonymous):

Solve 3x^2 - 4x - 6 = 0 by factoring.

OpenStudy (anonymous):

is the equation not, 3x^2-3x-6 ?

OpenStudy (anonymous):

3x^2 - 4x - 6 = 0

OpenStudy (anonymous):

the best way to do this is quadratic formula.

OpenStudy (anonymous):

I just get the numbers wrong... when I do it..

OpenStudy (anonymous):

is the sum from a textbook?

OpenStudy (dan815):

unfoil it

OpenStudy (dan815):

@Kainui

OpenStudy (anonymous):

I got x=4 x=6

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

I tried to quad formula...

OpenStudy (anonymous):

well the simplest way is the quad formula.

OpenStudy (anonymous):

\[\large ax^{2} + bx + c = 0\] \[\large 3x^{2} -4x -6 = 0\] a= 3 b = -4 c = -6

OpenStudy (anonymous):

The quadratic formula is \[\large \frac{ -b \pm \sqrt{b^{2} -4ac} }{ 2a }\] \[\large \frac{4 \pm \sqrt{4^{2} -4(3)(-6)} }{ 2(3) }\] \[\large = \frac{4 \pm \sqrt{16 +72} }{ 2(3) }\] \[\large = \frac{4 \pm \sqrt{88} }{ 2(3) }\]

OpenStudy (anonymous):

(3x-2)(x+3)

OpenStudy (anonymous):

\[\large = \frac{4 \pm 2\sqrt{22} }{ 2(3) }\] \[\large = \frac{2 \pm \sqrt{22} }{ (3) }\]

OpenStudy (anonymous):

set that equal to 0 and you get x=2/3 and x = -3

OpenStudy (anonymous):

(3x-2) (x+3) = 3x^2 -2x +9x -6 \[= 3x^{2} -7x +6\] We need -4x

OpenStudy (anonymous):

=3x^2 + 7x -6 ***